### v2

# Language-Specific Latent Process Hinders Cross-Lingual Performance

Abstract
Large language models (LLMs) are demonstrably capable of cross-lingual transfer, but can produce inconsistent output when prompted with the same queries written in different languages. To understand how language models are able to generalize knowledge from one language to the others, we apply the logit lens to interpret the implicit steps taken by LLMs to solve multilingual multi-choice reasoning questions. We find LLMs predict inconsistently and are less accurate because they rely on subspaces of individual languages, rather than working in a shared semantic space. While larger models are more multilingual, we show their hidden states are more likely to dissociate from the shared representation compared to smaller models, but are nevertheless more capable of retrieving knowledge embedded across different languages. Finally, we demonstrate that knowledge sharing can be modulated by steering the models’ latent processing towards the shared semantic space. We find reinforcing utilization of the shared space improves the models’ multilingual reasoning performance, as a result of more knowledge transfer from, and better output consistency with English.
1 Introduction
Humans have an innate ability to apply common knowledge and perform reasoning skills consistently across different languages, as the cognitive functions involved are largely independent of the languages they are learned in [5, 4]. Modern large language models (LLMs) are trained predominantly in English [24, 8] and Chinese [35]. To some degree, LLMs demonstrate cross-lingual capabilities [27], but the performance gap between languages remains [15, 13], and the models struggle to maintain consistent responses when they are probed by the same queries in different languages [20, 12, 7].
One possible explanation is that language models tend to develop language-specific neurons, rather than sharing them among multiple languages [29]. Nevertheless, language-specific neurons have never been shown to cause inconsistency in LLMs directly. Other work suggests that LLMs operate in a shared subspace closest to the English tokens, before projecting back to other languages in higher layers, for predictions in multilingual and multi-modal tasks [32, 34, 41]. These findings, however, fail to explain the lack of consistency across languages for the largest models [7] — otherwise, model outputs would converge toward English-prompted responses as multilingual performance improves.
This work investigates how language models generalize knowledge to multiple languages. We experiment on Gemma 2 and Qwen 2.5 models and show that:
Utilization of shared semantic space in multilingual processing facilitates knowledge transfer.
We first evaluate LLMs’ cross-lingual consistency and transfer on multi-choice queries. We find larger LLMs are indeed more accurate, but present large performance gaps, and remain inconsistent with poor cross-lingual transfer. We then investigate the intermediate steps taken implicitly by LLMs in solving the queries, and discover that models produce more divergent responses when operating in subspaces of different languages, rather than a shared internal representation that is closest to English hidden representation when processing multilingual prompts. We present evidence demonstrating a strong link between this phenomenon and the transfer performance of LLMs.
Implicit processing in English is more stable and accurate.
We interpret the hidden representations of LLMs using the logit lens [18, 32, 42, 34], which project intermediate hidden layers to the vocabulary space, to compute the latent distribution of answer choices. LLMs are capable of distinguishing multilingual answer choices in their hidden states, as indicated by our observation that the internal representations are closer to the embedding of the correct answers, relative to the other choices. Our results show the implicit process in English is more stable and more accurate than the process in the native languages of the prompts.
Nonetheless, larger models tend to reason in the native language of the prompt.
In line with previous work, we find models of smaller scale work predominantly in the English shared space. Larger models, however, dissociate their internal states from the dominant semantic space, and move towards the individual subspaces of the native languages. While we have shown this phenomenon to undermine multilingual reasoning performance, the impact is less pronounced on larger models. We attribute this to their capability of encoding and retrieving information in different languages independently with significantly more parameters.
Inducing processing in the shared semantic space improves cross-lingual transfer and model performance in reasoning tasks.
Building on current understanding of knowledge transfer, how can we improve cross-lingual consistency while retaining performance of individual languages? The final part of this work demonstrates that knowledge sharing and consistency can be modulated by reinforcing or diminishing processing in the English latent space. To this end, we introduce cross-lingual activation steering, inspired by prior work [30, 23]. We find adding a steering vector towards English produces more English-like predictions, leading to performance gains in smaller models; whereas adding an opposite vector results in more divergent output, with a significant dip in performance.
In summary, our work provides an account of the internal multilingual processing of LLMs and explains the limitations of current models in knowledge transfer. To further validate our hypotheses we explore cross-lingual activation steering. We recommend future work to develop more comprehensive methods that encode knowledge of different languages in the shared semantic space to maximize transfer between languages.
2 Methodology
2.1 Cross-lingual consistency and knowledge transfer
We define an LLM to be cross-lingually consistent when it produces the same output to the same queries written in different languages. Unlike Ifergan et al. [12], our measure of consistency is based upon multiple-choice question-answering tasks. To reduce language variability in the output space, we formulate the consistency between a language pair by their similarity in answer ranks, given by an LLM when prompted with the same questions and answer choices in their respective languages. Formally, for each of $i\in I$ questions in language $l$, $q_{i,l}\in Q_{l}$, paired with $j\in\{1..J\}$ answer choices $[a_{i,l}^{1},...,a_{i,l}^{J}]$, we rank the answer choices by their output probability $p(a_{i,l}^{j}|q_{i,l})$ given the question. We then combine the ranks $k(a_{i,l}^{j})$ of each answer within the set of choices:
$\mathbf{k}_{l}$ $=\langle k(a_{1,l}^{1}),...,k(a_{1,l}^{J}),k(a_{2,l}^{1}),...,k(a_{I,l}^{J})\rangle.$
Thus, the pairwise consistency between languages $l_{1}$ and $l_{2}$ is defined as the Spearman’s rank correlation coefficient between the rank vectors:
$\mathrm{consistency}({l_{1},l_{2}})=\rho(\mathbf{k}_{l_{1}},\mathbf{k}_{l_{2}}).$ (1)
Not all transfer between languages is beneficial—a model can be cross-lingually consistent but make many incorrect predictions. To approximate the rate of correct answers in a language propagates to another, we adapt our approach from Ifergan et al. [12] and estimate the positive transfer from $l_{1}$ to $l_{2}$ by the ratio of shared accurate responses to all accurate responses in $l_{1}$:
$\mathrm{tr}^{+}(l_{1},l_{2})=|F_{l_{1}}\cap F_{l_{2}}|/|F_{l_{1}}|.$ (2)
$F_{l}$ denotes the set of questions in $l$ that the model answers correctly. Conversely, $\neg F_{l}$ refers to the set of questions in $l$ where the model responses incorrectly, and $\neg F_{l\>\cap\>m}$ the questions in $l$ and $m$ where the model return the same incorrect responses. We compute the negative transfer from $l_{1}$ as:
$\mathrm{tr}^{-}(l_{1},l_{2})=|\neg F_{l_{1}\>\cap\>l_{2}}|/|\neg F_{l_{1}}|.$ (3)
We note that transfer rates defined in Equations 2 and 3 are directed and non-commutative, unlike consistency in Equation 1. Further, it is not possible to satisfy all the listed requirements without perfect accuracy—full consistency is not attainable without both maximum positive and negative transfer rates. Overall, the consistency, positive and negative transfer rates of a model can be computed by the expected values for all pairwise combinations of languages, $N=\{l_{1},l_{2}\in L\>|\>l_{1}\neq l_{2}\}$:
$\operatorname{\mathbb{E}}[\textrm{consistency}]$ $=\frac{1}{|N|}\sum_{l_{1},l_{2}\in N}\mathrm{consistency}({l_{1},l_{2}}),$ (4)
[…]
2.2 Interpreting the latent processes of multilingual tasks
We rely on the logit lens to interpret the latent representation of a transformer by applying its output softmax layer to the hidden states [18]. These intermediate output distributions have been found to provide insights and largely explain the final predictions of a model [32, 1, 34].
Logit lens is primarily designed to probe the latent distribution of the next token. To estimate the probability of a sequence in the latent space, we follow Zhong et al. [42] to allow past tokens to be included in the input, and decode from the hidden states to compute the final probabilities at each layer. Let $[x_{t},...,x_{T}]$ be the token indices of a phrase $\mathbf{w}$ following a query $\mathbf{q}$ with token indices $[x_{0},...,x_{t-1}]$, $h^{t}_{\ell}$ be the hidden state of layer $\ell$ at timestep $t$, and $O$ denote the embedding matrix. The latent probability of $\mathbf{w}$ at layer $\ell$, $p_{\ell}(\mathbf{w}|\mathbf{q})$, is equivalent to
$\prod_{\tau\in[t..T]}\mathrm{softmax}(Oh^{t}_{\ell})_{x_{\tau}}$ (7)
which we normalize by the length of $\mathbf{w}$: $T-t+1]{p_{\ell}(\mathbf{w}|\mathbf{q})}$.111 The approach is analogous to prior work that analyzes the latent representation of LLMs, where traces of reasoning [36] and abstract concepts [14] are found, providing insights into model predictions. The proposed method is also plausible because it is directly related to how an autoregressive model predicts the next token, where a textual answer can be sampled from the token distribution.
Based on these equations, we operationalize the relative similarity between the hidden states and embedding space in terms of the normalized probabilities of the answer choices, which are evaluated in their full textual forms given a query. We then rank the answer choices accordingly and compute the consistency and transfer between all language pairs, using the methods described in Section 2.1. For clarity, we use $\eta^{\ell}_{m}$ to denote ranked answer choices in language $m$ queried in language $\eta$, derived from intermediate layer $\ell$. For example, ${\textrm{jv}}^{2}_{\alpha}$ and ${\textrm{jv}}^{2}_{\textrm{en}}$ refer to the ranked output at the second layer in alphabetical labels, and English respectively, from a Javanese query, which is to be distinguished from ${\textrm{jv}}^{2}_{\textrm{jv}}$ where the ranked output is provided in Javanese.
Using the latent probabilities and the ranked output, we test if language models could reason in languages beyond English and extract information stored in different languages [32, 1, 41, 34, 12]. Based on the findings of Tang et al. [29] and Goldman et al. [7], we hypothesize that large-scale LLMs, with a greater number of language-specific neurons in the hidden layers, could induce higher activation and utilization of multilingual subspace rather than the shared semantic space. We also expect operating in individual language subspace rather than the shared subspace to be the cause of limited knowledge transfer and inconsistency across multilingual prompts.
2.3 Cross-lingual activation steering
To further test these hypotheses, we derive from model behavior steering literature [30, 23] and propose cross-lingual activation steering. Because the shared latent space is most similar to English [32], we expect pushing the latent processes towards English to induce processing in the shared space, allowing for knowledge sharing. On the other hand, pushing the process towards the original language is expected to cause more divergent predictions.
Following Rimsky et al. [23], we obtain the contrastive addition vector of multiple-choice questions for each languages, by computing the mean difference in activation as the model processes parallel queries in English and each of the languages, for each data set. The vector is extracted from the last token of the prompts, as the models prepare to generate an answer in the next token. Specifically, for a set of parallel queries in English and language $m$, $Q_{{\textrm{en},m}}$, the steering vector at layer $\ell$ from $m$ to English is computed as:
$v^{\ell}_{m\rightarrow\textrm{en}}=\frac{1}{|Q_{{\textrm{en},m}}|}\sum_{q_{\textrm{en}},q_{m}\in Q_{{\textrm{en},m}}}h_{\ell}(q_{\textrm{en}})-h_{\ell}(q_{m}),$ (8)
where $h_{\ell}(...)$ refers to the hidden activation at layer $\ell$ of the query’s last token. Like previous work [23], we use a scalar multiplier $\gamma$ to modulate the direction of change and the magnitude of the vectors. $\gamma>0$ induces latent process similar to being prompted in English, whereas $\gamma<0$ reinforces language-specific processing. At inference, the extracted addition vector is added to the hidden activation as $\gamma\cdot v^{\ell}_{m\rightarrow\textrm{en}}$ at the last token of the prompts.
3 Experimental setup
Models.
We experiment on Gemma 2 [24] and Qwen 2.5 [35] family of language models, which have been trained and instruction-tuned primarily in English. We use 2B, 9B and 27B variants of Gemma models and 3B, 7B and 32B of Qwen models. Note that the Qwen models have been tuned on multilingual instructions and are claimed to support 29 languages [22]. For all models, we compute the output distribution of the hidden states at an interval of four layers, ending at the final layer (e.g., layer 2, 6, …, 42; where 42 is the total number of layers in Gemma 2 9B).
Data.
We examine the latent process of the models using zero-shot prompts with three parallel multilingual tasks: Belebele for machine reading comprehension in 122 languages, created by human translators [2]; mMMLU for world knowledge and problem solving questions that have been manually translated and are available in 15 languages [9, 19]; and mHellaSwag for commonsense natural language inference in 10 languages, created with automatic translations [17, 39].
For multiple-choice questions, we associate each answer choice with an alphabetic characters (e.g., ‘A’), and evaluate the models based on the output distribution of the alphabets listed in the prompts. All prompt templates used are reported in Appendix A. We then derive based on Equation 7 the latent probabilities of the answer choices in i) their full textual form and ii) their parallel translations in English. We then follow Bandarkar et al. [2] to compute the accuracy of the models in these tasks.
4 Results
4.1 Cross-lingual performance and the impact of language relations
Table 1: Mean accuracy ($\uparrow$), consistency ($\uparrow$), positive ($\uparrow$) and negative ($\downarrow$) transfer. Larger models are more accurate, but the predictions are less consistent, with limited transfer across languages.
Belebele mMMLU mHellaSwag
acc. cons. $\textrm{tr}^{+}$ $\textrm{tr}^{-}$ acc. cons. $\textrm{tr}^{+}$ $\textrm{tr}^{-}$ acc. cons. $\textrm{tr}^{+}$ $\textrm{tr}^{-}$
Gemma 2 2B 47.1$\pm$13.8 .381 .623 .372 42.5$\pm$5.9 .540 .671 .509 44.8$\pm$8.3 .461 .656 .470
Gemma 2 9B 72.4$\pm$18.4 .427 .779 .273 59.0$\pm$8.2 .561 .771 .478 66.2$\pm$11.9 .558 .780 .418
Gemma 2 27B 75.1$\pm$18.2 .548 .803 .289 62.6$\pm$8.3 .639 .810 .521 69.9$\pm$12.3 .631 .807 .429
[…]
Qwen 2.5 7B 57.6$\pm$20.3 .383 .650 .278 54.9$\pm$10.6 .541 .724 .442 60.4$\pm$15.0 .474 .718 .381
Qwen 2.5 32B 70.3$\pm$20.9 .485 .757 .277 66.4$\pm$11.8 .567 .798 .443 70.0$\pm$14.2 .555 .796 .396
Figure 1: Positive transfer in all language pairs on mMMLU. Languages that are similar or are written in the same script are more likely to transfer with one another.
We report the accuracy and cross-lingual transfer performance of all models based on the ranked alphabet choices determined by the output distribution, i.e., between $x^{\lambda}_{\alpha}$ and $y^{\lambda}_{\alpha}$, $\lambda$ is the final layer, $x$ and $y$ are a language pair. Table 4.1 shows the mean accuracy across languages and the consistency and transfer rates (Equations 4, 5 and 6) for individual language models and benchmarks. Despite being more accurate, larger models such as Gemma 2 27B and Qwen 2.5 32B have more varied performances. In terms of consistency and transfer across languages, larger models offer little gain compared to their scale and accuracy, especially relative to the smaller models.
Figure 1 shows the pairwise positive transfer for Gemma 2 9B on mMMLU. Like Ifergan et al. [12], we find languages with known relations, such as Indic languages: Hindi (hi) and Bengali (bn); and Romance languages: Portuguese (pt), Spanish (es) and Italian (it); to be more closely connected and transfer more knowledge with each other. Indonesian (id) has higher rates of transfer with European languages, relative to Chinese (zh), Korean (ko) and Japanese (ja); indicating the role of shared scripts in knowledge transfer. The models are also less capable of generalizing predictions to and infer knowledge from low-resource languages: Swahili (sw) and Yoruba (yo). Knowledge learned in other languages remains inaccessible in yo as the models improve with increasing capacity.
4.2 Multilingual processing in the latent space
(a) Log ratio of latent probabilities of answer choices in the native language and English. Models assign more mass to the native language and operate in the language subspace when the log ratio exceeds zero.
(b) Mean accuracy in the latent space across languages. Correct answers can be partially identified in the hidden states. Chance-level accuracy is marked with a dashed line.
Figure 2: Mean log ratio of latent probabilities and latent accuracy across languages. Larger models tend to operate in a targeted subspace of individual languages, rather than the shared semantic space that is dominant in smaller models. English representation appears more stable and accurate. The shaded areas are 95% confidence intervals across languages.
To understand cross-lingual consistency and knowledge transfer, our next analyses explore the latent multilingual processing by computing the normalized probabilities of answer choices in their native form (nat.) and their English equivalents (en); based on Equation 7.
The log ratios between probability of the native and English answer choices in Figure 1(a) reveal that larger models tend to assign more mass to answers in the native form in the intermediate layers, as $\log p(\textrm{nat. choices})/p(\textrm{en choices})>0$. At layer 22 and beyond, the probability of the native language exceed that of English in Gemma 27B. We see a similar trend in Qwen 7B and 32B with English at layers 4 and 40 respectively. This indicates that large-scale language models operate in individual space that is closer to the token embeddings of the native language. Our results also suggest the contrary for smaller models: Gemma 2B/9B and Qwen 3B. Early layers of the models are similar to the embeddings of the native languages. The hidden states then shift to working predominantly in the English in the middle layers, where the ratio is below zero, before moving closer to the native language space again at the highest layers. This is in line with previous findings [29, 41] that most language-specific neurons are developed at the top and bottom layers of LLMs.
We further interpret the implicit steps taken across layers, by computing for all languages the accuracy of $x^{\ell}_{\textrm{en}}$ and $x^{\ell}_{x}$. In Figure 1(b), the latent predictions are shown to be noisy, however, exhibit accuracy above chance level in higher layers across data sets. Nonetheless, English latent predictions appear more accurate and stable in most cases.
[…]
Belebele mMMLU mHellaSwag
[…]
Gemma 2 2B $-0.6$∗∗∗ $-0.7$∗∗∗ $-0.8$∗∗∗ $-0.7$ $-0.9$∗∗∗ $-0.9$∗∗∗ $-0.2$ $-1.0$∗∗∗ $-1.0$∗∗∗
Gemma 2 9B $-0.5$∗∗∗ $-0.6$∗∗∗ $-0.6$∗∗∗ $-0.3$ $-0.5$ $-0.5$ $0.2$ $0.0$ $0.1$
Gemma 2 27B $-0.3$∗∗∗ $-0.1$ $-0.0$ $-0.6$ $-0.4$ $-0.4$ $-0.8$ $-0.9$∗∗ $-0.9$∗∗
[…]
Qwen 2.5 7B $-0.2$ $-0.2$ $-0.1$ $-0.7$∗∗ $-0.6$ $-0.6$ $-0.9$∗∗∗ $-0.9$∗∗∗ $-1.0$∗∗∗
Qwen 2.5 32B $-0.2$ $0.0$ $-0.0$ $-0.5$ $-0.4$ $-0.5$ $-0.9$∗∗ $-0.9$∗∗ $-0.9$∗∗
Taken together, our analyses imply that smaller models are limited to working within a shared representation, that is closest to English. Conversely, larger models transform the hidden states into subspace of individual languages, rather than the dominant representation. While it is possible for models to reason in a particular language, the process also hinders the models’ ability to infer from the other languages, contributing to cross-lingual inconsistency.
4.3 Modulating cross-lingual consistency with activation steering
(a) PCA projections of layer activations in Gemma 2 9B.
(b) PCA projections of layer activations in Gemma 2 27B.
Figure 3: Layer activations with 50 parallel questions from Belebele.
We have inferred in Section 4.2 that larger models process multilingual queries differently than smaller models. Figure 3 provides a visual comparison of the hidden activations of 50 parallel Belebele questions between Gemma 9B and 27B, where the dimension of the vectors is reduced with the projection of Principal Component Analysis (PCA) [33]. [^2] PCA is useful to assess if the hidden representations are linearly separable, which predetermines the efficacy of steering vectors [23]. While the activations can be grouped by languages in the early layers in Gemma 9B (Figure 2(a)), they are mostly aligned between layers 18 and 38, and are separated by their written scripts (Latin vs. non-Latin) in the final layer. We note however that activations of low-resource languages, such as Yoruba (blue squares), are the last to be mapped and continue to be disjoint from the other languages across hidden layers. [^3] We include analogous comparison between Qwen 3B and 32B in Appendix C.1.
In contrast, the latent space of Gemma 27B is more expansive, and the activations of the same languages are partially aligned with the others between layers 14 and 26, then become more tightly clustered throughout the layers. This confirms our previous analysis that larger models tend to operate in independent subspaces. Based on these observations, we expect steering vectors described in Section 2.3 to be most effective in the middle layers, where the hidden activations across languages better align, indicating the use of shared semantic space in the process.
Figure 4: Steered Gemma 2B. Note that $\gamma>0$ indicates an activation vector that pushes the model towards English processing in the latent space, and vice versa.
(a) Steered models ($\gamma=2.0$) evaluated on Belebele.
(b) Steered models ($\gamma=2.0$) evaluated on mMMLU.
Figure 5: Steering towards English is more effective on smaller models. Steering layers are shown at the top of the panels. The x-axes indicate the accuracy of non-steered models.
We extract steering vectors from 50 randomly sampled parallel questions from each data set, and evaluate the cross-lingual performance of the steered models on held-out data. Figure 4 reports a sweep of the steering vector multiplier, $\gamma$, on layers 6 and 10 of Gemma 2B. Positively steered models become increasingly accurate (peaking at $2\leq\gamma\leq 3$), are more consistent with English and receive more positive transfer from English, as the hidden states move towards English subspace. [^4] We observe generally consistent variability across $\gamma$, in terms of standard deviations and confidence intervals over languages. This is omitted in the graph to avoid visual clutter. Similar plots for Qwen models are included in Appendix C.2.
We also observe that steering vectors at the middle layers are the most impactful, because the hidden representations are intended to be aligned at this stage (Figure 3), to allow for language-agnostic task solving [41]. [^5] The performance of steered Gemma models across layers is reported in Figure 8 in Appendix C.3. Figure 5 shows the accuracy changes (%) in steered models across languages. In Figure 4(a), we see low-resource (accuracy) languages written in Latin script benefit the most from the method in Belebele; whereas the improvement in mMMLU (Figure 4(b)) is less notable without steering vectors specialized to each topics. [^6] Similar result can be observed in mHellaSwag, which is reported in Figure 9 in the Appendix. We also include corresponding analyses on positive transfer and consistency in Appendix C.4. The method is also less effective on larger models. We speculate on two possible reasons: i) the performance of larger models is saturated, with little headroom for improvement; ii) the hidden representations might become increasingly non-linear with scale, therefore linear perturbations with steering vectors are less useful on larger models.
5 Related work
Cross-lingual consistency and knowledge sharing.
Qi et al. [20] define pairwise cross-lingual knowledge consistency as the equality in ranked outputs scaled by softmax, whereas Ifergan et al. [12] and Goldman et al. [7] measure the extent to which a language pair are accurate in a similar way with factual questions. Our use of Spearman’s correlation in the evaluation of models’ consistency allows a more direct approach, while also taking into account the relative order of ranks. Across languages, script similarity has been proposed to be a major factor to drive knowledge transfer [20, 12, 7]. Our work also supports previous findings that models are capable of processing and storing information with independent representation of different languages [12], and explains previous findings that transfer performance increases with model size only up to 14B for Qwen 2.5, indicating the ability of larger models to retrieve knowledge learned from independent languages [7].
Cross-lingual representation alignment.
A growing body of work has shown that multilingual LLMs use English as a pivot language in the hidden layers [32], and share the subspace across languages and modalities [34]. Alabi et al. [1] find that the phenomenon also applies to models adapted to new languages, where finetuned adapters preserve most of the representation of the source language of the model, rather than working in a parallel isolated structure. Other works suggest the distinction between language-agnostic and language-specific neurons [41, 31], motivating a series of methods that improve multilingual representation and performance of LLMs with targeted supervision [40, 11]. The current study extends this understanding of multilingual models, and provides evidence for larger LLM’s capability to encode meaningful representation in languages beyond English.
English-centric multilingual reasoning.
Prior work has shown that explicit alignment with English reasoning improves multilingual performance. Existing approaches include English translations of multilingual queries [27, 43, 16], using English reasoning traces [27, 21, 37, 16], post-hoc translation from English answers [16], and trained alignment with English representation [38, 10, 26]. Our exploration on cross-lingual activation steering adds to this line of work, and offers a straightforward method to tap into English reasoning capability without training or additional inference costs.
6 Discussion
This work investigates the challenges of cross-lingual transfer and consistency in LLMs. We examine the implicit processing of multilingual queries and show LLMs can operate in a multilingual subspace. With more capacity, larger models exhibit a higher utilization of independent subspaces of individual languages, whereas smaller models are constrained to work predominantly in the shared semantic space. Models are also less likely to share knowledge when they derive multilingual predictions from subspaces of individual languages, rather than the shared semantic space. Our work also demonstrates that reinforcing processing in the shared space via activation engineering improves multilingual reasoning performance and knowledge transfer in smaller models.
Despite the contributions of this study, we address two limitations. For simplicity, we adopt logit lens to evaluate answer choices in phrases. Other alternatives allow for more expressive interpretations, but require model training [3, 6]. While there is noticeable gain in performance with steered models, the method is only tested on multi-choice questions, and requires extraction of specific activation vectors for individual languages, which may not generalize to other use cases. We also caution against the use of English-centric methods beyond reasoning tasks, which might reinforce cultural and linguistic biases in LLMs [28, 25, 37]. It is however possible for future approaches to learn more stable internal representations when processing multilingual queries, which allow for complete retrieval of knowledge learned in different languages.
Acknowledgments
Special thanks to Grace Chung, Honglin Yu, Alan Ansell and Fan Jiang for their insights and research support.
Appendix A Prompt template
Table 3: Prompt templates used for Belebele, mMMMLU and mHellaSwag.
Belebele
Given the following question and answer choices, output the letter corresponding to the correct answer. <passage> Question: <question> A: <mc answer 1> B: <mc answer 2> C: <mc answer 3> D: <mc answer 4> Answer:
mMMLU
Given the following question and answer choices, output the letter corresponding to the correct answer. Question: <question> A: <option a> B: <option b> C: <option c> D: <option d> Answer:
mHellaSwag
Given the following question and answer choices, output the letter corresponding to the correct answer. Question: <ctx> A: <ending 1> B: <ending 2> C: <ending 3> D: <ending 4> Answer:
Table 3 lists the prompt templates used in our experiments reported in the main text.
Appendix B Compute resources
We conducted experiments on models with fewer than 10B parameters using NVIDIA A100 GPUs (80GB) and on larger models using H100 GPUs. All models were able to fit on a single GPU for inference, and the experiments required approximately 200 hours of compute time.
Appendix C Additional results on cross-lingual activation steering
C.1 Hidden activations in Qwen models
Figure 6 provides a comparison of hidden activations between Qwen 3B and 32B models. Like Gemma, the hidden representation in the middle layers is more aligned in the smaller model than the larger model.
(a) PCA projections on activations in Qwen 2.5 3B.
(b) PCA projections on activations in Qwen 2.5 32B.
Figure 6: Activations in Qwen models on 50 parallel questions from Belebele, analogous to Figure 3 in the main text.
C.2 Varying multiplier on steered Qwen models
We perform a sweep of the steering vector multiplier, $\gamma$, on Qwen 3B and 32B, and report the results in Figure 7.
(a) Qwen 2.5 3B steered at layer 12.
(b) Qwen 2.5 32B steered at layer 12.
Figure 7: Qwen 3B steered towards latent English processing is more accurate, more consistent with English and receive more positive transfer from English. The experiment on Qwen 32B with mMMLU and mHellaSwag, however, reveal potential improved consistency at the expense of accuracy, particularly for models with large enough capacity to store information independently from English.
C.3 Steering Gemma models across layers
Steering is most effective in the middle layers, as evidenced by Figure 8.
Figure 8: Gemma 2 models steered across layers, where $\gamma\in\{2.0,-2.0\}$. Shaded areas are 95% confidence intervals over languages.
C.4 Accuracy on mHellaSwag, positive transfer and consistency changes in steered models.
Steered results on mHellaSwag is reported in Figure 9.
Figure 9: Accuracy changes in mHellaSwag. The improvement is more substantial in smaller models.
We show the changes in positive transfer from English and consistency with English in Figures 10 and 11 respectively. Overall, the results demonstrate that cross-lingual steering is effective on smaller models; whereas it is more difficult to induce stable and consistent transfer in larger models with simple activation addition.
(a) Belebele
(b) mMMLU
(c) mHellaSwag
Figure 10: Changes in positive transfer from English in steered models ($\gamma=2.0$) across tasks.
(a) Belebele
(b) mMMLU
(c) mHellaSwag
Figure 11: Changes in consistency with English in steered models ($\gamma=2.0$) across tasks.